OpenStudy (anonymous):

how you solve for y? lol x=(2y+3)/(3-2y)

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

well u have \[ \large x=\frac{2y+3}{3-2y} \] what is dividing on one side changes sides to multiply so \[ \large x(3-2y)=2y+3 \]

OpenStudy (helder_edwin):

now distribute \[ \large 3x-2xy=2y+3 \] \[ \large 3x-3=2y+2xy=y(2+2x) \] therefore \[ \large y=\frac{3x-3}{2+2x} \]

OpenStudy (anonymous):

nice but in 2y+2xy can't you factor 2y? why did you factor the y only :D

OpenStudy (helder_edwin):

because u r solving for y. u could factor 3 in the numerator and 2 in the denominator but u gain nothing.

OpenStudy (anonymous):

oooh thank you very much :d